# Weighted Orlicz *-algebras on locally elliptic groups

### v1

@@ 8.1. Locally elliptic groups with the bounded index property
    If $q>1$ and $p\in(1,\infty)$ such that $11/p+1/q=$ 1, then $\omega_{f,q}$ is an $L^{p}$-weight.
@@ 8.2. Contraction groups
If we further assume that the contraction group $G$ is totally disconnected, then we may choose $U\leqslant G$ a compact open subgroup, and it follows from $\alpha$ being contractive that $G=\bigcup_{n=1}^{\infty}\alpha^{-n}(U)$. Thus, $G$ is locally elliptic since $\alpha^{n}(U)$ is compact for each $n$, and $G$ has the bounded index property since the indices $[\alpha^{-n-1}(U):\alpha^{-n}(U)]$ are constant with $n$ varying. We state this below.
@@
is a non-solvable torsion contraction group when equipped with the right-shift automorphism as its contractive automorphism. This group, however, is also Hermitian since it is an $[FC]^{-}$-group [34, Theorem 2.6]. But torsion contraction groups need not be $[FC]^{-}$-groups either, with the Heisenberg group over $\mathbb{F}_{p}(\!(t)\!)$ providing an example of a torsion contraction group that is not an $[FC]^{-}$-group.
@@ 8.3. Unipotent linear algebraic group over non-archimedean local fields
Let $k$ be a non-archimedean local field and define $U_{n}(k)$ to be the group of all upper-triangular $n$-dimensional unipotent matrices over $k$. Given $x\in k$ with absolute value $<1$, one checks that conjugation by the diagonal matrix $\text{diag}(1,x,x^{2},\dots,x^{n-1})$ is a contractive automorphism of $U_{n}(k)$. Thus, it follows that $U_{n}(k)$ is a locally elliptic group with the bounded index property by the discussion in the previous section, and since these properties pass to closed subgroups, we get the following result.
@@ 8.5. Horocyclic groups of automorphism of trees and scale groups
As far as the author is aware, it is even an open question as to whether pright-normal-factor-semidirect-$\mathbb{Q}_{p}\rtimes\mathbb{Z}$ is Hermitian (c.f. [48, Section 3.6]).

### v2

@@ 8.1. Locally elliptic groups with the bounded index property
    If $q>1$ and $p\in(1,\infty)$ such that $1/p+1/q=1$, then $\omega_{f,q}$ is an $L^{p}$-weight.
@@ 8.2. Contraction groups
If we further assume that the contraction group $G$ is totally disconnected, then we may choose $U\leqslant G$ a compact open subgroup, and it follows from $\alpha$ being contractive that $G=\bigcup_{n=1}^{\infty}\alpha^{-n}(U)$. Thus, $G$ is locally elliptic since $\alpha^{n}(U)$ is compact for each $n$, and $G$ has the bounded index property since the indices $\alpha^{-n-1}(U):\alpha^{-n}(U)]$ are constant with $n$ varying. We state this below.
@@
is a non-solvable torsion contraction group when equipped with the right-shift automorphism as its contractive automorphism. This group, however, is also Hermitian since it is an $FC]^{-}-$group [34, Theorem 2.6]. But torsion contraction groups need not be $FC]^{-}-$groups either, with the Heisenberg group over $\mathbb{F}_{p}(\!(t)\!)$ providing an example of a torsion contraction group that is not an $FC]^{-}-$group.
@@ 8.3. Unipotent linear algebraic group over non-archimedean local fields
Let $k$ be a non-archimedean local field and define $U_{n}(k)$ to be the group of all upper-triangular $n$-dimensional unipotent matrices over $k$. Given $x\in k$ with absolute value <1<1, one checks that conjugation by the diagonal matrix $\text{diag}(1,x,x^{2},\dots,x^{n-1})$ is a contractive automorphism of $U_{n}(k)$. Thus, it follows that $U_{n}(k)$ is a locally elliptic group with the bounded index property by the discussion in the previous section, and since these properties pass to closed subgroups, we get the following result.
@@ 8.5. Horocyclic groups of automorphism of trees and scale groups
As far as the author is aware, it is even an open question as to whether $\mathbb{Q}_{p}\rtimes\mathbb{Z}$ is Hermitian (c.f. [48, Section 3.6]).